solve by substitution 2x+6y=14
-3x+y=29 What is the soultion set ( -3,2) am I right

Answer by Alan3354(69443)

2x+6y=14 --> 2*(-3) + 6*2 = 14 That's not right.
-3x+y=29
----------
What is the soultion set ( -3,2) am I right
No. How did you get that?
Show your work via the Thank You note.